OpenStudy (anonymous):

what is the solution to -15=-4m+5?

OpenStudy (anonymous):

Are you trying to find the solution to the whole equation or just the variable?

OpenStudy (anonymous):

the whole equation

OpenStudy (anonymous):

\[-15=-4m+5\] Swapping m to the left \[-4m + 5 = -15\] \[-4m+5-5=-15-5\] \[-4m=-20\] \[\frac{ -4m }{ 4 } = \frac{ -20 }{ -4 }\] \[m=5\]

OpenStudy (anonymous):

\[5+-4\times5=-15\]